# Hybrid-Emba3D: Geometry-Aware and Cross-Path Feature Hybrid Enhanced State Space Model for Point Cloud Classification

### v1

@@ 3.2 Local geometric pooling with geometry-feature coupling mechanism
Figure 4: Illustration of LGP. This process selects neighboring points through FPS and KNN, extracts normalized features from the center points, and element-wise multiplication with the broadcast extended local geometric weights to achieve the fusion of features and geometric information, followed by a shared MLP for advanced feature extraction.
According to Section 3.1 and Figure 4, $L$ local domains with $k$ neighbors are constructed using KNN around each center point. The center point of these local domains is ${{P}_{\text{center}}}\in{{\mathbb{R}}^{L\times 3}}$ , and the coordinates of the neighborhood point are ${{k}_{i}}$.
For the local feature normalization, given the center point feature and its K nearest neighbor feature ${{\tilde{F}}_{K}}\in{{\mathbb{R}}^{L\times K\times C}}$, the relative feature offset $\Delta F={{F}_{K}}-{{F}_{C}}$ is first calculated and channel normalized:
${{\tilde{F}}_{K}}=\frac{\Delta F}{\sqrt{Var(\Delta F)+\varepsilon}}$ (13)
The normalized neighbor features and the center point features are concatenated along the channel, and the propagation features can be generated through learning an affine transformation:
@@
Where: $\gamma$ and $\beta$ are trainable proportional coefficients and shift vectors, respectively.
That implements feature-aware adaptive updates by explicitly encoding the center-neighborhood feature relationship.
For the local geometric patch normalization, firstly compute the mean coordinate $\mu$ of $M$ points within the patch, then convert the coordinates of each point into offsets $\Delta p={{p}_{j}}-u$ relative to the patch centroid to eliminate global translational effects. Subsequently, calculate the coordinate standard deviation $\sigma={\sqrt{\sum\nolimits_{x=1}^{M}{||\Delta{{p}_{j}}||_{2}^{2}+\varepsilon}}}/{M}\;$, and normalize the coordinates by $p_{j}^{{}^{\prime}}={\Delta{{p}_{j}}}/{\sigma}\;$ to eliminate local scale differences. This process ensures that the geometric encoding achieves translation-scale invariance. The process can be formalized as:
$p_{j}^{{}^{\prime}}=\frac{{{p}_{j}}-\mu}{\sqrt{\frac{1}{M}\sum\nolimits_{k=1}^{M}{||{{p}_{k}}-\mu||_{\text{2}}^{\text{2}}+\varepsilon}}}$ (15)
Where: ${{p}_{j}},{{p}_{k}}\in{{\mathbb{R}}^{3}}$ are 3D coordinates of the $j$-th or $k$-th point in the local geometric patch; $\mu$ is the Centroid of the geometric patch, $\mu={\sum\nolimits_{k=1}^{N}{{{p}_{k}}}}/{M}\;$; $j$, $k$ are point indices, where $j$ is used for calculating the offset of individual points, and $k$ is used for summation traversal over all points; $\varepsilon$ is the numerical stability term.
@@
$\Delta{{p}_{i}}={{p}_{i}}-{{p}_{c}},$ (16)
${{w}_{i}}=\exp(-||\Delta{{p}_{i}}|{{|}_{2}}),$ (17)
$W=Broadcast({{w}_{i}})\in{{\mathbb{R}}^{2C\times L\times K}}$ (18)
${{F}_{weighted}}={{{\hat{F}}}_{k}}\odot W$ (19)
Where: ${{p}_{i}},{{p}_{c}}\in{{\mathbb{R}}^{3}}$ represent the coordinates of patch points and the center point; ${{w}_{i}}\in{{\mathbb{R}}^{L\times K\times 1}}$ is a weight coefficient based on distance attenuation, ${{\hat{F}}_{k}}$ is a normalized feature tensor, with a shape of $2C$×$L$×$K$; $\odot$ represents element-wise multiplication; $Broadcast(\cdot)$ extends the weight vector to align with the feature dimension.
The final feature ${{F}_{agg}}\in{{\mathbb{R}}^{2C\times L}}$ is obtained using the Softmax weighted average to achieve adaptive aggregation of neighborhood features:
${{F}_{agg}}=\frac{\sum\nolimits_{k=1}^{K}{{{F}_{weighted}}\centerdot\exp({{F}_{weighted}})}}{\sum\nolimits_{k=1}^{K}{\exp({{F}_{weighted}})}}$ (20)
Finally, channel alignment is completed through the shared MLP.
@@
3.3 Bidirectional SSM with the Collaborative feature enhancer
@@
Figure 5: Details of Collaborative feature enhancer
Given the input tensor $X\in{{\mathbb{R}}^{B\times C\times L}}$ divided into $g$ groups along the channel dimension:
$X^{\prime}=\operatorname{Reshape}(X,[B\cdot g,\frac{C}{g},L])$ (21)
@@
$\phi({{X}_{i}})={{F}_{soft}}({{P}_{avg}}({{X}_{i}}))\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (24)
Where: ${F}_{soft}$ is the row-wise Softmax.
@@
$\text{output}^{\prime}=X^{\prime}\odot W\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (28)
$\text{output}=\operatorname{Reshape}(\text{output}^{\prime},(B,C,L))$ (29)
CoFE employs dual pathways to capture local-global contexts, where a dynamic weight matrix explicitly fuses cross-scale features via bidirectional interaction (${X}_{1}\rightleftarrows{X}_{2}$). This interaction computes spatial affinities to eliminate unidirectional bias and adaptively models non-uniform dependencies in long sequences, delivering a lightweight attention enhancement for BiSSM (see Figure 6).
@@
The unidirectional nature of Mamba restricts global dependency capture in point cloud processing. Unlike Vision Mamba’s[53] token-level horizontal flipping, a reverse path through vertical flipping of feature channels is constructed. The resulting BiSSM module is formulated as:
$F^{\prime}=\text{CoFE}(Con{{v}_{1\times 1}}(F))$ (30)
[…]
@@
This approach eliminates reliance on point cloud token order, enabling the model to focus on inherent distribution patterns of feature vectors rather than artificially constructed sequential orders. While Mamba’s core mechanism captures long-range dependencies through parameterized state transition equations, it demonstrates insufficient sensitivity to local transient features. The CoFE component complements the SSM framework by directly modeling inter-position relationships. The ${X}_{2}$ pathway reduces the limitations of SSMs in capturing local details through localized context aggregation. Subsequent ${X}_{1}\rightleftarrows{X}_{2}$ interaction facilitates complementary integration of feature correlation and provides more discriminative input to SSMs, enabling SSMs to dynamically perceive global contextual patterns while effectively modeling complex sequential dependencies.
@@ 4.2 Object classification
Training Settings: For fine-tuning, the pre-training employs the Point-MAE framework with masked self-supervised learning on the ShapeNetCore[3] dataset. Unless explicitly specified, consistent training parameters are maintained between scratch training and fine-tuning. The Hybrid-Emba3D classification task header utilizes a 3-layer MLP classifier (256→256→$C$ channels). In the ModelNet40 experiment, adopting the standard division (9843 train/2468 test), sampling 1024 points, and using scale&translation augmentation, while ScanObjectNN followed the original train/test split, sampling 2048 points, and using the rotation augmentation. The experimental results are shown in Table 1.
Table 1: Experimental results of ModelNet40 and ScanObjectiNN. Compared with methods of different architectures: 3D understanding architectures, Transformer-based architectures, •Diffusion-driven frameworks, Mamba-based architectures. * denotes the application of voting strategy. The results of other models are sourced from publicly published papers.
@@
Hybrid-Emba3D Point-MAE 95.58+1.58 93.80+0.27 93.12+0.87 93.09+2.04 16.96 3.95
Hybrid-Emba3D* Point-MAE 95.99+1.99 95.18+1.38 93.80+1.55 94.34+3.29 16.96 3.95
Comparison with existing SOTA methods: Hybrid-Emba3D achieves SOTA performance across multiple benchmarks. Without pretraining, the model attains 94.0% accuracy on ModelNet40, surpassing the comparable Mamba3D by 0.6%, while achieving the highest OBJ_BG accuracy of 93.63%. When employing Point-MAE pretraining, Hybrid-Emba3D breaks the 95.58% accuracy barrier on ModelNet40, further reaching 95.99% with the voting strategy, establishing a new record for this benchmark. Hybrid-Emba3D outperforms SOTA Transformer counterparts (PointGPT-S, Point-FMAE) by 1.99% and 0.84% in accuracy, while reducing Parameters by 41.9% and 39.1%, respectively. Compared to similar Mamba-based models (Pointmamba, Mamba3D), it achieves 0.89% and 1.0% accuracy gains on ModelNet40 and ScanObjectNN-PB_T50_RS (the most challenging benchmark) with only exchanged 0.06M Parameters and 0.05G FLOPs.
Feature visualization:High-dimensional features are reduced via t-SNE to a 2D projection distribution to visualize the clustering effect of features. As shown in Figure 7. On ModelNet40, the feature distribution is relatively scattered, and there are ambiguous inter-class boundaries, indicating that the Hybrid Emba3D needs to improve its ability to distinguish specific categories. While OBJ_BG and PB_T50_RS demonstrate prominent clustering tendencies with minor inter-class overlaps, OBJ_ONLY achieves superior class separation. The model maintains robust clustering performance under PB_T50_RS interference scenarios, confirming its resilience to environmental disturbances.
Figure 7: Details of Collaborative feature enhancer
Figure 8: Details of Collaborative feature enhancer
@@
Comparing OBJ_ONLY and OBJ_BG, the presence of background increases the misjudgment rate of locally structurally similar objects. For example, the accuracy of category-0 (bag) decreases from 100% to 88%, and the confusion rate of category-5 (chair) and 9 (shelf) reaches 13%. Conversely, it may also reduce the misclassifications rate caused by feature overlap, such as the accuracy rate of category-10 and 12 increased from 86% and 75% to 95% and 88%, respectively, and the misjudgment rate of category-9 (shelf) and 10 (table), category-11 (bed) and category-12 (pillow) drops from 9% and 12% to 0%. Under interference (PB_T50_RS), the misclassification distribution has shifted, meaning misjudged samples are redistributed among error categories while maintaining a stable overall misjudgment rate.
@@
LGP CoFE-BiSSM ModelNet40 1k P ScanObjectNN P (M) F (G)
@@
Hybrid-Emba3D 93.3±5.0 98.9±1.3 90.4±5.7 95.6±3.9
4.3 Ablation study
To validate the complementary nature and distinct performance boundaries of core components LGP and CoFEBiSSM, across diverse scenarios, we conduct ablation studies on the Hybrid-Emba3D benchmarked on 3D point cloud classification tasks. The results as shown in Table 2
LGP: LGP is a core component focusing on local feature extraction. This component captures subtle geometric variations and topological structures on object surfaces by constructing non-linear transformation mechanisms within local neighborhoods. When solely activated, LGP achieves 94.00% accuracy on ModelNet40, demonstrating its strong geometric feature representation in ideal, background-free scenarios. Furthermore, it achieves accuracy gains of +0.97% (OBJ_BG), +1.04% (OBJ_ONLY), and +0.35% (PB_T50_RS) compared to the deactivated baseline, further validating its capability to resist background clutter and minor deformations through local geometric modeling. It is commendable that it introduces zero additional parameters (0M) and computational costs (0G FLOPs), highlighting superior lightweight design advantages.
CoFE-BiSSM: CoFE-BiSSM enhances long-range dependencies through multi-scale feature fusion and bidirectional modeling. It achieves 92.05% accuracy on OBJ_BG, slightly below LGP’s 92.57%, but with excellent efficiency (0.03M Parameters/0.09G FLOPs). In the ModelNet40 (93.56%) and OBJ-ONLY (90.02%) scenarios without background interference, the performance is relatively limited, while the rotation disturbance sensitivity on PB_T50_RS (90.32%) reveals insufficient local geometric modeling. Experiments demonstrate its effectiveness in global context reasoning for complex background tasks, but it requires integration with local feature extraction modules to address geometric sensitivity.
Joint effect: Component collaboration yields significant performance gains. At the peak of 94.00% (+0.61%) in ModelNet40, all subsets of ScanObjectNN surpassed the single component configuration comprehensively, with OBJ-BG improving by 2.03%, OBJ-ONLY improving by 2.41%, and PB_T50_S improving by 0.73%. They achieve feature complementarity with minimal overhead. In summary, LGP and CoFE-BiSSM focus on local geometric and global multi-scale modeling, respectively. While their strengths are scenario-specific, joint deployment enables complementary features and enhances generalization performance on different datasets.
4.4 Few-shot Learning
Results: Without pretraining, Hybrid-Emba3D achieves 90.5% on 5-Way 10-Shot, surpassing Transformer (87.8±5.2) but lagging behind DGCNN-CrossPoint (92.5%). Scaling to 20-shot boosts accuracy by 6% (96.0%). For 10-Way tasks, it outperforms DGCNN-CrossPoint. With self-supervised pretraining, it is worth noting that Hybrid Emba3D outperforms PointGPT-S (98.6%) in 5-Way 20 Shot, achieving a quasi SOTA performance of 98.9%. However, the advantage has not been generalized to other settings. In the 5/10-Way task, the accuracy rate increased by 5.6% and 5.2% when the sample size increased from 10-Shot to 20-Shot, and the standard deviation dropped sharply by 74% and 32%, indicating that the model is sensitive to data sparsity, and data enhancement can significantly improve performance and stability.

### v2

@@ 3.2 Geometry-Feature Coupled Pooling
Figure 4: Illustration of GFCP. Local neighborhoods are constructed based on KNN, and normalized features and coordinates of neighborhood points are extracted. These neighborhood features are concatenated with the central feature and processed through the affine transformation to generate the augmented feature. Concurrently, geometric weights are computed based on neighborhood coordinates and broadcast on the channel dimension. The features are aggregated after weighted fusion. Finally, advanced features are extracted through the shared MLP.
According to Section 3.1 and Figure 4, $G$ local domains are constructed around each center point. The central features and coordinates of these local domains are ${{F}_{c}}\in{{\mathbb{R}}^{G\times C}}$ and ${{P}_{c}}\in{{\mathbb{R}}^{G\times 3}}$. The $k$ nearest neighbor points are calculated via Euclidean distance within local regions to establish geometric neighborhoods capturing point cloud spatial topology, with neighborhood features ${{F}_{k}}\in{{\mathbb{R}}^{G\times k\times C}}$ and coordinates ${{P}_{k}}\in{{\mathbb{R}}^{G\times k\times 3}}$.
For the local feature extension, the relative feature offset $\Delta F={{F}_{K}}-{{F}_{C}}$ is first calculated between neighboring features and central features, and channel normalized:
${{\tilde{F}}_{K}}=\frac{\Delta F}{\sqrt{Std(\Delta F)+\varepsilon}}$ (9)
Where: ${{\tilde{F}}_{k}}\in{{\mathbb{R}}^{G\times k\times C}}$ is the normalized neighborhood features.
The normalized neighborhood features and the center point features are concatenated along the channel, and the Extended features can be generated through learning an affine transformation:
@@
Where: $\gamma$ and $\beta$ are the trainable scaling coefficient and bias vector for the channel dimension, respectively.
That implements dynamically adjusts the feature distribution and channel importance by explicitly encoding the center-neighborhood feature relationship and adaptive parameters $\gamma$, $\beta$, thereby achieving feature-aware adaptive updates.
For local geometric weighting, the generation of geometric weights needs to be based on normalized neighborhood coordinates. Firstly, convert the coordinates of the neighborhood point into offsets relative $\Delta P={{P}_{k}}-{{P}_{c}}$ to the central coordinate to eliminate global translational effects. Subsequently, calculate the neighborhood coordinate standard deviation, and normalize the neighborhood coordinates to eliminate local scale differences. The process can be formalized as:
${{\tilde{P}}_{k}}=\frac{\Delta P}{\sqrt{Std(\Delta P)+\varepsilon}}$ (11)
Where: ${{\tilde{P}}_{k}}\in{{\mathbb{R}}^{G\times k\times 3}}$ is the normalized neighborhood coordinates.
@@
${{w}_{i}}=\exp(-||{{{\tilde{P}}}_{k}}|{{|}_{2}}),$ (12)
${{W}_{geo}}=Broadcast({{w}_{i}})\in{{\mathbb{R}}^{G\times k\times 2C}}$ (13)
Where: ${{w}_{i}}\in{{\mathbb{R}}^{L\times K\times 1}}$ is a weight coefficient based on distance attenuation to reduce the impact of distant points on extended features; $Broadcast(\cdot)$ expands the weight vector to align with the feature dimension, where all channels within the same neighborhood share identical weights, achieving differentiated weighting exclusively in the spatial dimension.
Finally, performs element-wise multiplication between geometric weights and expanded features, thereby coupling local features with geometric structural information.
${{F}_{weighted}}={{\hat{F}}_{k}}\odot{{W}_{geo}}$ (14)
Where: ${{F}_{weighted}}\in{{\mathbb{R}}^{G\times k\times 2C}}$ is the geometrically weighted feature; $\odot$ represents element-wise multiplication;
The final local feature ${{F}_{agg}}\in{{\mathbb{R}}^{G\times 2C}}$ is obtained using the nonlinear Softmax-like Pooling to achieve adaptive aggregation of neighborhood features:
${{F}_{agg}}=\frac{\sum\nolimits_{k=1}^{K}{({{F}_{weight}}\odot\exp({{F}_{weight}}))}}{\sum\nolimits_{k=1}^{K}{\exp({{F}_{weight}})}}$ (15)
[…]
@@
3.3 BiSSM with the Collaborative feature enhancer
Figure 5: Details of Collaborative feature enhancer
@@ 3.3 BiSSM with the Collaborative feature enhancer
Given the input feature $X\in{{\mathbb{R}}^{B\times C\times L}}$ divided into $g$ groups along the channel dimension:
$X^{\prime}={Reshape}(X,[B\cdot g,\frac{C}{g},L])$ (16)
@@
$\phi({{X}_{i}})={{F}_{soft}}({{P}_{avg}}({{X}_{i}}))\in{{\mathbb{R}}^{(B\cdot g)\times 1\times{}^{C}\!\!\diagup\!\!{}_{g}\;}}$ (19)
Where: ${F}_{soft}$ is the Softmax along a specific dimension.
@@
${output}^{\prime}=X^{\prime}\odot W\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (23)
${output}={Reshape}({output}^{\prime},(B,C,L))$ (24)
CoFE employs dual pathways to capture local-global contexts, where a dynamic weight matrix explicitly fuses cross-scale features via bidirectional interaction (${X}_{1}\rightleftarrows{X}_{2}$). This interaction computes spatial affinities to eliminate unidirectional bias and adaptively models non-uniform dependencies in long sequences, delivering a lightweight attention enhancement for SSM (see Figure 6).
@@
The unidirectional nature of Mamba restricts global dependency capture in point cloud processing. Unlike Vision Mamba’s [42] token-level horizontal flipping, Mamba3D [12] constructs a reverse path through vertical flipping of feature channels. The resulting CoFE-BiSSM module is formulated as:
$F^{\prime}={Conv}_{1\times 1}({CoFE}(F))$ (25)
${CoFE-BiSSM}=Linear(F^{\prime}+Fli{{p}_{C}}(F^{\prime}))$ (26)
@@
This approach eliminates reliance on point cloud token order, enabling the model to focus on inherent distribution patterns of feature vectors rather than artificially constructed sequential orders. While Mamba’s core mechanism captures long-range dependencies through state transition equations’ parameter dynamization, it demonstrates insufficient sensitivity to local transient features. The CoFE component complements the SSMs framework by directly modeling inter-position relationships. The ${X}_{2}$ pathway reduces the limitations of SSMs in capturing local details through localized context aggregation. Subsequent ${X}_{1}\rightleftarrows{X}_{2}$ interaction facilitates complementary integration of feature correlation and provides more discriminative input to SSMs, enhancing BiSSM to dynamically perceive global contextual patterns while effectively modeling complex sequential dependencies.
@@ 4.2 Object classification
Training Settings: For fine-tuning, the pre-training employs the Point-MAE framework with masked self-supervised learning on the ShapeNetCore [45] dataset. Unless explicitly specified, consistent training parameters are maintained between scratch training and fine-tuning. The HyMamba classification task header utilizes a 3-layer MLP classifier (256→256→$C$ channels). In the ModelNet40 experiment, adopting the standard division (9843 train/2468 test), sampling 1024 points, and using scale&translation augmentation, while ScanObjectNN followed the original train/test split, sampling 2048 points, and using the rotation augmentation. The experimental results are shown in Table 1.
Table 1: Experimental results of ModelNet40 and ScanObjectiNN.
@@
E-Mamba [13] NeuCom 25 Point-MAE 94.7 94.32 92.94 91.98 13.78 7.49
HyMamba Point-MAE 95.58+1.58 93.80+0.27 93.12+0.87 93.09+2.04 16.96 3.95
HyMamba* Point-MAE 95.99+1.99 95.18+1.38 93.80+1.55 94.34+3.29 16.96 3.95
  • 1
    Compared with methods of different architectures: 3D understanding architectures, Transformer-based architectures, •Diffusion-driven frameworks, Mamba-based architectures. * denotes the application of voting strategy.
Comparison with existing SOTA methods: HyMamba achieves SOTA performance across multiple benchmarks. Without pretraining, the model attains 94.0% accuracy on ModelNet40, surpassing the comparable Mamba3D by 0.6%, while achieving the highest OBJ_BG accuracy of 93.63%. When employing Point-MAE pretraining, HyMamba breaks the 95.58% accuracy barrier on ModelNet40, further reaching 95.99% with the voting strategy, establishing a new record for this benchmark. HyMamba outperforms SOTA Transformer counterparts (PointGPT-S, Point-FMAE) by 1.99% and 0.84% in accuracy, while reducing Parameters by 41.9% and 39.1%, respectively. Compared to similar Mamba-based models (Pointmamba, Mamba3D), it achieves 0.89% and 1.0% accuracy gains on ModelNet40 and ScanObjectNN-PB_T50_RS (the most challenging benchmark) with only exchanged 0.06M Parameters and 0.05G FLOPs.
Notably, HyMamba demonstrates clear advantages over the latest E-Mamba in both accuracy (+0.88% on ModelNet40, +1.11% on PB_T50_RS) and computational efficiency (3.95G vs 7.49G FLOPs) when using pretrained weights. Although E-Mamba introduces geometric information in the process of reordering, it is fundamentally different from the explicit and direct geometric feature coupling of HyMamba. With pre-trained weights from ShapeNetCore providing a rich initial representation, this distinction becomes critical. GFCP can leverage this rich foundation to more accurately simulate complex local geometry, ultimately unlocking the full potential of the model to achieve significant performance improvements.
Feature visualization: High-dimensional features are reduced via t-SNE to a 2D projection distribution to visualize the clustering effect of features. As shown in Figure 7. On ModelNet40, the feature distribution is relatively scattered, and there are ambiguous inter-class boundaries, indicating that the HyMamba needs to improve its ability to distinguish specific categories. While OBJ_BG and PB_T50_RS demonstrate prominent clustering tendencies with minor inter-class overlaps, OBJ_ONLY achieves superior class separation. The model maintains robust clustering performance under PB_T50_RS interference scenarios, confirming its resilience to environmental disturbances.
Figure 7: t-SNE diagram of HyMamba on the ModelNet40 and ScanobjectNN.
@@
Figure 8: Confusion matrix of HyMamba on the ModelNet40 and ScanobjectNN.
Comparing OBJ_ONLY and OBJ_BG, the presence of background increases the misjudgment rate of locally structurally similar objects. For example, the accuracy of category-0 (bag) decreases from 100% to 88%, and the confusion rate of category-5 (chair) and 9 (shelves) reaches 13%. Conversely, it may also reduce the misclassifications rate caused by feature overlap, such as the accuracy rate of category-10 and 12 increased from 86% and 75% to 95% and 88%, respectively, and the misjudgment rate of category-9 (shelves) and 10 (table), category-11 (bed) and category-12 (pillow) drops from 9% and 12% to 0%. Under interference (PB_T50_RS), the misclassification distribution has shifted, meaning misjudged samples are redistributed among error categories while maintaining a stable overall misjudgment rate.
4.3 Ablation study
To validate the complementary nature and distinct performance boundaries of core components GFCP and CoFEBiSSM, across diverse scenarios, we conduct ablation studies on the HyMamba benchmarked on 3D point cloud classification tasks. The results as shown in Table 2
@@ 4.3 Ablation study
GFCP CoFE-BiSSM ModelNet40 1k P ScanObjectNN P (M) F (G)
@@
GFCP: GFCP is a core component focusing on local feature extraction. This component captures subtle geometric variations and topological structures on object surfaces by constructing non-linear transformation mechanisms within local neighborhoods. When solely activated, GFCP achieves 94.00% accuracy on ModelNet40, demonstrating its strong geometric feature representation in ideal, background-free scenarios. Furthermore, it achieves accuracy gains of +0.97% (OBJ_BG), +1.04% (OBJ_ONLY), and +0.35% (PB_T50_RS) compared to the deactivated baseline, further validating its capability to resist background clutter and minor deformations through local geometric modeling. It is commendable that it introduces zero additional parameters (0M) and computational costs (0G FLOPs), highlighting superior lightweight design advantages.
CoFE-BiSSM: CoFE-BiSSM enhances long-range dependencies through multi-scale feature fusion and bidirectional modeling. It achieves 92.05% accuracy on OBJ_BG, slightly below GFCP’s 92.57%, but with excellent efficiency (0.03M Parameters/0.09G FLOPs). In the ModelNet40 (93.56%) and OBJ-ONLY (90.02%) scenarios without background interference, the performance is relatively limited, while the rotation disturbance sensitivity on PB_T50_RS (90.32%) reveals insufficient local geometric modeling. Experiments demonstrate its effectiveness in global context reasoning for complex background tasks, but it requires integration with local feature extraction modules to address geometric sensitivity.
Joint effect: Component collaboration yields significant performance gains. At the peak of 94.00% (+0.61%) in ModelNet40, all subsets of ScanObjectNN surpassed the single component configuration comprehensively, with OBJ-BG improving by 2.03%, OBJ-ONLY improving by 2.41%, and PB_T50_S improving by 0.73%. They achieve feature complementarity with minimal overhead. In summary, GFCP and CoFE-BiSSM focus on local geometric and global multi-scale modeling, respectively. While their strengths are scenario-specific, joint deployment enables complementary features and enhances generalization performance on different datasets.
4.4 Few-shot Learning
@@
HyMamba 93.3±5.0 98.9±1.3 90.4±5.7 95.6±3.9
Results: Without pretraining, HyMamba achieves 90.5% on 5-Way 10-Shot, surpassing Transformer (87.8±5.2) but lagging behind DGCNN-CrossPoint (92.5%). Scaling to 20-shot boosts accuracy by 6% (96.0%). For 10-Way tasks, it outperforms DGCNN-CrossPoint. With self-supervised pretraining, it is worth noting that HyMamba outperforms PointGPT-S (98.6%) in 5-Way 20 Shot, achieving a quasi SOTA performance of 98.9%. However, the advantage has not been generalized to other settings. In the 5/10-Way task, the accuracy rate increased by 5.6% and 5.2% when the sample size increased from 10-Shot to 20-Shot, and the standard deviation dropped sharply by 74% and 32%, indicating that the model is sensitive to data sparsity, and data enhancement can significantly improve performance and stability.